OpenStudy (anonymous):

what are the y-intercepts of y=x^2-2x-24?

OpenStudy (anonymous):

Set x=0, to find y intercept y = -24

OpenStudy (anonymous):

mii choices was (-8,6) (-6,4) (6,-4) (-8,3) i went with (-6,4) is that right

OpenStudy (amistre64):

none of them are good since none of them have x equal to zero

OpenStudy (amistre64):

and why would you go with (-6,4) ??

OpenStudy (anonymous):

y=-24 is the ans

OpenStudy (anonymous):

laura what made u write thar

OpenStudy (anonymous):

(6, -4)

OpenStudy (amistre64):

perhaps its typoed and meant: what are the "x"-intercepts of y=x^2-2x-24? (x-6)(x+4) = 0 (6,-4) would be correct for that then, otherwise ...... nada